# Does hot QCD have a conformal manifold in the chiral limit?

### 10 Mar 2026 v1

@@ Abstract
Early lattice simulations seemed to agree with these predictions [4, 5, 6, 7, 8, 9, 10, 11], but recent lattice simulations [12, 13, 14, 15, 16, 17, 18, 19] attribute the earlier $N_{f}\geq 3$ results to discretization artifacts and find no evidence of a first-order transition for any $N_{f}\geq 2$ at $\mu_{B}=0$ or imaginary $\mu_{B}$ [19]. Lattice results thus suggest that the chiral transition is second-order, or at most weakly first-order, for any $N_{f}\geq 2$ at $\mu_{B}=0$ and along a line extending to imaginary $\mu_{B}$.
@@
$\rho_{B}(x)\equiv\frac{i}{\beta}\int_{0}^{\beta}\!\!\!\mathrm{d}\tau\,j_{B}^{0}(x,\tau)$ (8)
@@
The quantum picture is convenient for describing 3D infrared (IR) effective theories at length scales $\ell\gg\beta$. These theories are unitary 3D quantum field theories (QFTs) for any real $\theta_{B}$, and $\rho_{B}$ flows to a hermitian operator. When the 3D effective theory becomes scale-invariant at a critical point, unitarity is expected to enhance scale invariance to full conformal invariance; see e.g. Refs. [51, 52, 53, 54, 55, 56].
@@
$\begin{gathered}\!\!\!\int\!\!\mathcal{D}\phi\exp\!\left\{\!-\frac{1}{2g^{2}}\!\!\int\!\!\mathrm{d}^{3}\!x\,\mathrm{tr}\partial_{a}\phi\,\partial^{a}\!\phi^{\dagger}\!+i\!\!\int\!\!\theta_{B}\frac{\mathrm{tr}(\phi^{\dagger}\mathrm{d}\phi)^{3}}{24\pi^{2}}\!\right\}\!.\end{gathered}$ (17)
@@
In this high-$T$ gapped phase, continuous chiral symmetry flows to an emergent $2$-form $U(1)$ symmetry [70] and the 3D anomaly (9) is matched via symmetry transmutation [71]; we explain this high-$T$ IR effective theory in the Supplementary Material S1.
In summary, chiral symmetry is broken in the low-$T$ phase and restored in the high-$T$ phase, while the discrete symmetry at $\theta_{B}=\pi$ is broken [^2] If a CFT matching the symmetry and the anomaly existed, a second-order transition at $\theta_{B}=\pi$ without symmetry breaking would have a chance to occur at relatively lower $T$ in the high-$T$ phase [67, 25, 26, 27]; see the Supplemental Material S2 for some comments. in the high-$T$ phase and restored in the low-$T$ phase. This complementarity of symmetry realizations is a fingerprint of mixed anomalies.
Consistent phase diagrams. The low-$T$ and high-$T$ phases are separated by a critical line $\theta_{B}\mapsto T_{c}(\theta_{B})$. Motivated by recent lattice results [12, 13, 14, 15, 16, 17, 18, 19], we postulate that this line generically consists of natural second-order points. The IR physics of a natural critical point is a 3D CFT whose neutral sector under the symmetry (15) has precisely one relevant primary scalar operator $\mathcal{O}_{E}$, to which the $T$ perturbation flows. The operator $\rho_{B}$ also flows to $\mathcal{O}_{E}$ at $\theta_{B}\neq 0,\pi$, or at $\theta_{B}=0,\pi$ if spontaneous symmetry breaking (18) occurs. Such 3D CFTs should be unitary by our earlier argument.
@@
If the changes in the IR physics are abrupt, a neighborhood of $\theta_{B}=0$ on the critical line only cares about the breaking $G\to H$, and should fit the Landau paradigm, i.e. be described by a natural fixed point of the Ginzburg-Landau model of Ref. [3]. Minimally, it suffices to have one phase transition at $\theta_{B}=\pi$ and a Ginzburg-Landau description for the rest of the critical line. If the transition at $\theta_{B}=\pi$ is first-order (Landau scenario, Fig. 1 left), discrete symmetry breaking (18) with $\langle\rho_{B}\rangle\neq 0$ occurs at $\theta_{B}=\pi$. There must then be a window at $\theta_{B}=\pi$ where chiral and discrete symmetries are broken, with an Ising transition at its lower edge. If instead the transition at $\theta_{B}=\pi$ is second order (Landau-DQCP scenario, Fig. 1 middle), it must be a beyond-Landau critical point analogous to condensed-matter DQCPs [29, 30, 31]. The operator $\rho_{B}$ flows to a relevant (or marginally relevant) operator at this DQCP.
@@
where $\Phi(x)$ is an $N_{f}$-by-$N_{f}$ complex matrix field, with the symmetry of  (15) acting on $\Phi(x)$ in the same way as $\phi(x)$. The question is whether this model hits a natural critical point as we vary $\mu$. [^4] One may speculate that the model (22) can flow to a fixed point with emergent $U(1)_{A}$ symmetry that has the $\lambda_{3}$ term as an irrelevant deformation [3, 72, 73, 74, 75, 76, 77, 78]. However, as we discuss in the Supplemental Materials S3, at present no convincing evidence supports this possibility in the model (22).
@@
At mean-field level, the $\lambda_{3}$ term drives the transition first-order for $N_{f}=3$, but allows an $O(2N_{f}^{2})$ second-order transition at $\mu=0$ for $N_{f}\geq 4$. Unfortunately, beyond mean-field level, extensive searches using various methods have found no stable fixed points in the model (22) for all $N_{f}\geq 3$ [3, 79, 80, 81, 82, 83, 20, 22]. We thus consider the Landau and Landau-DQCP scenarios to be disfavored for $N_{f}\geq 3$.
@@
to the $SU(N_{f})_{\theta_{B}}$ CFT. This $SU(N_{f})_{\theta_{B}}$ field theory must describe the 3D IR physics of finite-$T$ massless QCD for all $T$. It should flow to a gapped phase for, say, $\mu>0$, and a Nambu-Goldstone gapless phase for $\mu<0$.
@@
To see whether the $(2+\epsilon)$D UV fixed point extrapolates to $\epsilon=1$, we study the scaling dimension of the lowest-dimension primary scalar $\mathcal{O}_{\phi}$ in the $\boldsymbol{N_{f}}\otimes\overline{\boldsymbol{N_{f}}}$ representation using the three-loop results of Refs. [96, 97, 98]:
$\Delta_{\phi}\sim\left(1-\frac{1}{N_{f}^{2}}\right)\left[\epsilon-\frac{1}{2}\epsilon^{2}+\frac{1}{2}\epsilon^{3}\right]+\mathcal{O}\left(\epsilon^{4}\right)\,.$ (24)
To estimate $\Delta_{\phi}$ in 3D, we apply Padé-Conformal-Borel (PCB) resummation [99, 100, 101, 102, 103, 104]; see Ref. [105] for a review. Errors are estimated by varying the Padè order and dominant Borel-plane singularity location; see the Supplemental Material S4. We summarize our 3D estimates for small $N_{f}$ in Table 1, which do satisfy the unitarity bound $\Delta_{\phi}\geq\frac{1}{2}$ in 3D.
$N_{f}$ 22 33 44
$\Delta_{\phi}$ $0.53\pm 0.03$ $0.63\pm 0.03$ $0.67\pm 0.04$
Table 1: Estimates for $\Delta_{\phi}$ in the 3D $SU(N_{f})_{0}$ CFT obtained from the Padé-Conformal-Borel (PCB) resummation of the three-loop $(2+\epsilon)$D expansion.
Conformal bootstrap [106] studies exist for 3D CFTs with symmetry $SU(N_{f})_{L}\times SU(N_{f})_{R}]/\mathbb{Z}_{N_{f}}$. For $N_{f}=2$, Ref. [107] found an island with $0.517\lesssim\Delta_{\phi}\lesssim 0.521$. For $N_{f}=3$, Ref. [108] found a rough kink at $\Delta_{\phi}=0.62\pm 0.01$ in one fusion channel. We see preliminary agreement at $N_{f}=2$ and $N_{f}=3$ between our $(2+\epsilon)$D estimate of $\Delta_{\phi}$ and the bootstrap estimates.
@@
Outlook. Motivated by recent lattice results [12, 13, 14, 15, 16, 17, 18, 19] on the QCD chiral transition, we argued that only three basic scenarios (Fig. 1) are consistent with a natural second-order critical line. Lattice QCD simulations should distinguish them by measuring the leading IR behaviors of $\rho_{B}$ at $\theta_{B}=0,\pi$ on the critical line, as summarized in Table 2.
@@
The conformal-manifold scenario is the only option not disfavored for $N_{f}\geq 3$ by current evidence. However, it requires the $SU(N_{f})_{\theta_{B}}$ universality class, whose existence is an open question of independent interest. Demonstrating it will require improved conformal bootstrap bounds on 3D CFTs with the symmetry (15) and further development of the $(2+\epsilon)$D expansion. Lattice Monte Carlo simulations of the 3D $SU(N_{f})$ PCM using Lüscher’s admissibility method [111, 112, 113, 114] (see also Ref. [115]) should be valuable.
While we assume a second-order critical line, a weakly first-order transition is not excluded by Refs. [12, 13, 14, 15, 16, 17, 18, 19]. If such transitions are controlled by nearby non-unitary complex CFTs [116, 117], such as the $(2+\epsilon)$D fixed points, much of our discussion would carry over.
@@
Acknowledgments. We are very grateful to Thomas Dumitrescu, Amey Gaikwad, Evgenii Ievlev, David B. Kaplan, Joseph Kapusta, Zohar Komargodski, Jiaxin Qiao, Laurence Yaffe, and Yuan Xin for very helpful discussions. OpenAI’s Codex 5.4, Anthropic’s Claude Opus 4.6, and Refine.ink were helpful in proofreading the manuscript. A.C. and S.C. are supported by Simons Foundation award number 994302, as part of the Simons Collaboration on Color Confinement and QCD Strings. R.D.P. is supported by the U.S. Department of Energy under contract DE-SC0012704, and by the Alexander von Humboldt Foundation.
@@
One may speculate that such an appropriate 3D CFT can simply be provided by colorless free fermions, because the 3D anomaly (9) can be matched by an $N_{f}$-flavor free massless 4D Dirac fermions at finite temperatures. At $\theta_{B}=\pi$, one Matsubara mode becomes massless, leading to the 3D effective theory at $\ell\gg\beta$:
@@
Nevertheless, as already pointed out by Refs. [25, 27], although it matches the anomaly (9) of QCD, the free-fermion CFT (S2.1) does not match the symmetry (1) of QCD. A colorless fermionic object in the fundamental representation of $SU(N_{f})_{L}$ or in the anti-fundamental representation of $SU(N_{f})_{R}$ can exist in QCD only when
@@
When this happen, this colorless fermionic object necessarily carries a baryon number $q_{B}$ satisfying $q_{B}=1\mod N_{f}$. For example, with physical $N_{c}=3$, the above conditions are satisfied by $N_{f}=2$, but are violated by $N_{f}=3$. Therefore, a gapless window at $\theta_{B}=\pi$ described by the free-fermion CFT (S2.1) is at least conceivable for $N_{f}=2$, but is immediately excluded for $N_{f}=3$. Of course, a more complicated free-fermion CFT might simultaneously match the symmetry and the anomaly, and we leave a detailed exploration of such possibilities to the future.
@@
The three-loop beta function of the $(2+\epsilon)$D $SU(N_{f})$ PCM is given by [97, 98] (in the convention of Ref. [96])
@@
$\beta(g^{2})\ \sim\ \epsilon g^{2}-\frac{N_{f}}{4}g^{4}-\frac{N_{f}^{2}}{32}g^{6}-\frac{3N_{f}^{3}}{512}g^{8}+\mathcal{O}\left(g^{10}\right)\,.$ (S4.1)
The UV fixed point is thus at
$g^{2}_{*}\ \sim\ \frac{4\epsilon}{N_{f}}\left(1-\frac{1}{2}\epsilon+\frac{1}{8}\epsilon^{2}\right)+\mathcal{O}\left(\epsilon^{4}\right)\,.$ (S4.2)
@@
The three-loop anomalous dimension of $\phi_{ij}$, the vertex operator corresponding to the matrix elements in the $\boldsymbol{N_{f}}\otimes\overline{\boldsymbol{N_{f}}}$ representation of $SU(N_{f})_{L}\times SU(N_{f})_{R}$, is given by [96]
$\gamma_{\phi}(g^{2})\ \sim\ \frac{\left(N_{f}^{2}-1\right)}{4N_{f}}\left(g^{2}+\frac{3N_{f}^{2}}{128}g^{6}\right)+\mathcal{O}\left(g^{8}\right)\,.$ (S4.3)
@@
Note that a vertex operator has the natural engineering dimension 0, even in $(2+\epsilon)$D. Namely, our $\gamma_{\phi}$ is the analogy to $\frac{1}{2}\zeta$ in the $O(N)$ nonlinear sigma model of Ref. [137, Section 4.1.3] (see especially their Footnote 35). As a result, the scaling dimension of $\phi_{ij}$ at the UV fixed point is given by
$\Delta_{\phi}(\epsilon)\ \equiv\ \gamma_{\phi}(g_{*}^{2})\ \sim\ \left(1-\frac{1}{N_{f}^{2}}\right)\left[\epsilon-\frac{1}{2}\epsilon^{2}+\frac{1}{2}\epsilon^{3}\right]+\mathcal{O}\left(\epsilon^{4}\right)\,.$ (S4.4)
The fact that the $N_{f}$-dependence factorizes at this order in perturbation theory is likely to be an accident, and should not hold generically.
Our goal here is to extrapolate $\Delta_{\phi}(\epsilon)$ to $\epsilon=1$. The most naive extrapolation would neglect $\mathcal{O}(\epsilon^{4})$ terms and find
$\Delta_{\phi}(1)=1-\frac{1}{N_{f}^{2}}$ (S4.5)
This cannot be remotely reliable because of the terrible truncation artifacts of alternating series at $\epsilon=1$, as well as the fact that underlying perturbative series should be asymptotic. To get more reasonable estimates of $\Delta_{\phi}$ in 33D, we need a sophisticated resummation method.
Here we extrapolate from $(2+\epsilon)$D to 33D using the Padé-Conformal-Borel (PCB) resummation [99, 100, 101, 102, 103, 104], see also Ref. [105] for a review. Given a truncated series with $K$ terms for $K\gg 1$, PCB achieves fractional errors scaling as $K^{-4}$, significantly better than the $K^{-2}$ scaling of standard Padé-Borel summation. Here $K=3$, so we cannot expect to be in the asymptotic regime where PCB is fully reliable. Nevertheless, we expect it to be significantly better than a naive extrapolation.
For convenience, let us introduce
$F(\epsilon)\ \equiv\ \frac{\Delta_{\phi}(\epsilon)}{\epsilon\left(1-\frac{1}{N_{f}^{2}}\right)}\ \sim\ 1-\frac{1}{2}\epsilon+\frac{1}{2}\epsilon^{2}+\mathcal{O}(\epsilon^{3})$ (S4.6)
Our goal is to construct an approximation of the actual function $F(\epsilon)$ and examine its behavior at $\epsilon=1$. The first step in this resummation scheme is to pass to the Borel plane by replacing $\epsilon^{n}$ with $\delta^{n}/n!$ in the series, to obtain
$\mathcal{F}(\delta)\ \sim\ 1-\frac{\delta}{2}+\frac{\delta^{2}}{4}+\mathcal{O}\left(\delta^{3}\right)\,.$ (S4.7)
Since this series seems to be alternating, the nearest Borel-plane singularity is likely on the negative real axis, but we cannot reliably estimate its location due to the truncation. Let us therefore assume that the nearest Borel-plane singularity is at $\delta=-p$ with $p>0$, and then treat $p$ as a free parameter that we will vary to get an estimate of the truncation error in our resummation.
In Padé-Borel resummations, one usually constructs a Padé approximant to the Borel-plane series, and then integrates it against the Borel kernel $\frac{1}{\epsilon}e^{-\delta/\epsilon}$ over the domain $0,\infty)$. The idea of the PCB resummation is to improve the convergence of this procedure by first performing a conformal transformation $\delta\mapsto z(\delta)$ given by
$z(\delta)=\frac{\sqrt{1+\frac{\delta}{p}}-1}{\sqrt{1+\frac{\delta}{p}}+1}.$ (S4.8)
with the inverse $\delta(z)=\frac{4pz}{(1-z)^{2}}$. This conformal transformation maps the branch cut $\infty+i\varepsilon\to-p\to-\infty-i\varepsilon$ to the arc $1+i\varepsilon\to-1\to 1-i\varepsilon$ on the unit circle. It also maps the Borel integral domain $0,\infty)$ to $[0,1)$, which then lies within the convergence disk of the conformal-transformed Borel-transformed series
$\overline{\mathcal{F}}(z)\ \equiv\ \mathcal{F}(\delta(z))\ \sim\ 1-2pz+4p\left(p-1\right)z^{2}+\mathcal{O}\left(z^{3}\right).$ (S4.9)

### 14 Apr 2026 v2

@@ Abstract
Early lattice simulations seemed to agree with these predictions [16, 73, 66, 5, 72, 83, 33, 67], but recent lattice simulations [29, 35, 91, 133, 134, 135, 76, 30, 75] attribute the earlier $N_{f}\geq 3$ results to discretization artifacts and find no evidence of a first-order transition for any $N_{f}\geq 2$ at $\mu_{B}=0$ or imaginary $\mu_{B}$ [30]. Lattice results thus suggest that the chiral transition is second-order, or at most weakly first-order, from $N_{f}=2$ up to the onset of the conformal window[75], at $\mu_{B}=0$ and along a line extending to imaginary $\mu_{B}$.
@@
$\rho_{B}(x)\equiv\frac{i}{\beta}\int_{0}^{\beta}\!\!\!\mathrm{d}\tau\,j_{B}^{0}(x,\tau)\,$, (8)
@@
The quantum picture is convenient for describing 3D infrared (IR) effective theories at length scales $\ell\gg\beta$. These theories are unitary 3D quantum field theories (QFTs) for any real $\theta_{B}$, and $\rho_{B}$ flows to a hermitian operator thereof. When the 3D effective theory becomes scale-invariant at a critical point, unitarity is expected to enhance scale invariance to full conformal invariance; see e.g. Refs. [101, 38, 93, 34, 95, 88].
@@
$\begin{gathered}\!\!\!\int\!\!\mathcal{D}\phi\exp\!\left\{\!-\frac{1}{g^{2}}\!\!\int\!\!\mathrm{d}^{3}\!x\,\mathrm{tr}\partial_{a}\phi\,\partial^{a}\!\phi^{\dagger}\!+i\!\!\int\!\!\theta_{B}\frac{\mathrm{tr}(\phi^{\dagger}\mathrm{d}\phi)^{3}}{24\pi^{2}}\!\right\}\!.\end{gathered}$ (17)
@@
In this high-$T$ gapped phase, continuous chiral symmetry flows to an emergent $2$-form $U(1)$ symmetry [49] and the 3D anomaly (9) is matched via symmetry transmutation [107]; we explain this high-$T$ IR effective theory in the Supplemental Material S1.
In summary, chiral symmetry is broken in the low-$T$ phase and restored in the high-$T$ phase, while the discrete symmetry at $\theta_{B}=\pi$ is broken [^2] If a CFT matching the symmetry and anomaly existed, a second-order transition at $\theta_{B}=\pi$ without symmetry breaking could occur at lower $T$ in the high-$T$ phase [127, 131, 94, 77]; see the Supplemental Material S2 for comments. in the high-$T$ phase and restored in the low-$T$ phase. This complementarity of symmetry realizations is a fingerprint of mixed anomalies.
Consistent phase diagrams. The low-$T$ and high-$T$ phases are separated by a critical line $\theta_{B}\mapsto T_{c}(\theta_{B})$. Motivated by recent lattice results [29, 35, 91, 133, 134, 135, 76, 30, 75], we postulate that this line generically consists of natural second-order points with emergent scale invariance. The IR physics of such a natural critical point is a 3D unitary CFT whose neutral sector under the symmetry (15) has precisely one relevant primary scalar operator $\mathcal{O}_{E}$, to which the $T$ perturbation flows. The operator $\rho_{B}$ also flows to $\mathcal{O}_{E}$ at $\theta_{B}\neq 0,\pi$, or at $\theta_{B}=0,\pi$ if spontaneous symmetry breaking (18) occurs.
@@
If the changes in the IR physics are abrupt, a neighborhood of $\theta_{B}=0$ on the critical line only cares about the breaking $G\to H$, and should fit the Landau paradigm, i.e. be described by a natural fixed point of the Ginzburg-Landau model of Ref. [99]. Minimally, it suffices to have one phase transition at $\theta_{B}=\pi$ and a Ginzburg-Landau description for the rest of the critical line. If the transition at $\theta_{B}=\pi$ is first-order (Landau scenario, Fig. 1 left), discrete symmetry breaking (18) with $\langle\rho_{B}\rangle\neq 0$ occurs at $\theta_{B}=\pi$. There must then be a window at $\theta_{B}=\pi$ where chiral and discrete symmetries are broken, with an Ising transition at its lower edge. [^3] Such a window is disfavored by current lattice evidence [14, 28]. If instead the transition at $\theta_{B}=\pi$ is second order (Landau-DQCP scenario, Fig. 1 middle), it must be a beyond-Landau critical point analogous to condensed-matter DQCPs [110, 109, 111]. The operator $\rho_{B}$ flows to a relevant (or marginally relevant) operator at this DQCP.
@@
where $\Phi(x)$ is an $N_{f}$-by-$N_{f}$ complex matrix field, with the symmetry (15) acting on $\Phi(x)$ in the same way as $\phi(x)$. The question is whether this model hits a natural critical point as $\mu$ varies. [^5] One may speculate that the model (22) flows to a fixed point with emergent $U(1)_{A}$ symmetry, with the $\lambda_{3}$ term irrelevant [99, 23, 82, 22, 44, 6, 96, 27]. However, as discussed in the Supplemental Material S3, no convincing evidence supports this possibility in the model (22).
@@
At mean-field level, the $\lambda_{3}$ term drives the transition first-order for $N_{f}=3$, but allows an $O(2N_{f}^{2})$ second-order transition at $\mu=0$ for $N_{f}\geq 4$. Unfortunately, beyond mean-field level, extensive searches have found no stable fixed points in the model (22) for all $N_{f}\geq 3$ [99, 18, 19, 2, 40, 102, 41, 39]. We thus consider the Landau and Landau-DQCP scenarios to be disfavored for $N_{f}\geq 3$.
@@
to the $SU(N_{f})_{\theta_{B}}$ CFT. This $SU(N_{f})_{\theta_{B}}$ field theory must describe the 3D IR physics of finite-$T$ massless QCD for all $T$. It should flow to a gapped phase for $\mu>0$ and a Nambu-Goldstone phase for $\mu<0$.
@@
To see whether the $(2+\epsilon)$D UV fixed point extrapolates to $\epsilon=1$, we study the scaling dimensions of $\mathcal{O}_{E}$ and the lowest-dimension primary scalar $\mathcal{O}_{\phi}$ in the $\boldsymbol{N_{f}}\otimes\overline{\boldsymbol{N_{f}}}$ representation using the three-loop results of Refs. [87, 62, 63] and the four-loop results of Ref. [119]:
$\Delta_{E}$ $=2-\frac{1}{2}\epsilon^{2}-\frac{1}{4}\epsilon^{3}-\left(\frac{5}{16}+\frac{9\zeta_{3}}{4N_{f}^{2}}\right)\epsilon^{4}+\mathcal{O}(\epsilon^{5}),$ (24a)
$\Delta_{\phi}$ $=\left(1-\frac{1}{N_{f}^{2}}\right)\left[\epsilon-\frac{1}{2}\epsilon^{2}+\frac{1}{2}\epsilon^{3}\right]+\mathcal{O}\left(\epsilon^{4}\right).$ (24b)
We estimate $\Delta_{E}$ and $\Delta_{\phi}$ in the 3D $SU(N_{f})_{0}$ CFT using Padé-Borel resummation; see the Supplemental Material S4. We summarize our 3D estimates with small $N_{f}$ in Table 1, which are consistent with unitarity $\Delta_{E},\Delta_{\phi}\geq\frac{1}{2}$ and relevance $\Delta_{E},\Delta_{\phi}<3$. More terms in the $(2+\epsilon)$D expansion are needed to reduce the extrapolation errors, which are significant for $\Delta_{E}$.
$N_{f}$ 22 33 44 55
$\Delta_{E}$ $1.94\pm 0.56$ $1.76\pm 0.67$ $1.64\pm 0.71$ $1.56\pm 0.72$
$\Delta_{\phi}$ $0.54\pm 0.02$ $0.64\pm 0.02$ $0.67\pm 0.02$ $0.69\pm 0.02$
Table 1: Estimates for $\Delta_{E}$ and $\Delta_{\phi}$ in the 3D $SU(N_{f})_{0}$ CFT obtained from Padé-Borel resummations of the known $(2+\epsilon)$D expansions.
Conformal bootstrap [100] studies exist for 3D CFTs with symmetry $SU(N_{f})_{L}\times SU(N_{f})_{R}]/\mathbb{Z}_{N_{f}}$. For $N_{f}=2$, Ref. [79] found an island with $1.64\lesssim\Delta_{E}\lesssim 1.68$ and $0.517\lesssim\Delta_{\phi}\lesssim 0.521$. For $N_{f}=3$, Ref. [80] found a rough kink at $\Delta_{\phi}=0.62\pm 0.01$. We see preliminary agreement at $N_{f}=2$ and $N_{f}=3$ between our $(2+\epsilon)$D estimates and the bootstrap estimates.
@@
Outlook. Recent lattice results [29, 35, 91, 133, 134, 135, 76, 30, 75] on the QCD chiral transition motivate our conclusion that only three scenarios (Fig. 1) are consistent with a natural second-order critical line. Lattice QCD simulations should distinguish them by measuring the leading IR behaviors of $\rho_{B}$ at $\theta_{B}=0,\pi$ on the critical line, in Table 2.
@@
The conformal-manifold scenario is the only option not disfavored for $N_{f}\geq 3$ by current evidence. However, it requires the $SU(N_{f})_{\theta_{B}}$ universality class, whose existence is an open question of independent interest. Demonstrating it will require improved conformal bootstrap bounds on 3D CFTs with the symmetry (15) and further development of the $(2+\epsilon)$D expansion. Lattice Monte Carlo simulations of the 3D $SU(N_{f})$ PCM using Lüscher’s admissibility method [86, 106, 114, 85] (see also Ref. [20]) might be valuable.
While we assume a second-order critical line, a weakly first-order transition is not excluded by Refs. [29, 35, 91, 133, 134, 135, 76, 30, 75]. If such transitions are controlled by nearby non-unitary complex CFTs [54, 53], such as the $(2+\epsilon)$D fixed points, much of our discussion would carry over.
@@
Acknowledgments. We are very grateful to Thomas Dumitrescu, Amey Gaikwad, Evgenii Ievlev, Theodore Jacobson, David B. Kaplan, Joseph Kapusta, Zohar Komargodski, Owe Philipsen, Jiaxin Qiao, Srimoyee Sen, Laurence Yaffe, and Yuan Xin for very helpful discussions. OpenAI’s Codex 5.4, Anthropic’s Claude Opus 4.6, and Refine.ink were helpful in proofreading the manuscript. A.C. and S.C. are supported by Simons Foundation award number 994302, as part of the Simons Collaboration on Confinement and QCD Strings. R.D.P. is supported by the U.S. Department of Energy under contract DE-SC0012704, and by the Alexander von Humboldt Foundation.
@@
One may speculate that such an appropriate 3D CFT can simply be provided by colorless free fermions, because the 3D anomaly (9) can be matched by an $N_{f}$-flavor free massless 4D Dirac fermion at finite temperatures. At $\theta_{B}=\pi$, one Matsubara mode becomes massless, leading to the 3D effective theory at $\ell\gg\beta$:
@@
Nevertheless, as already pointed out by Ref. [77], although it matches the anomaly (9) of QCD, the free-fermion CFT (S2.1) does not match the symmetry (1) of QCD. A colorless fermionic object in the fundamental representation of $SU(N_{f})_{L}$ or in the anti-fundamental representation of $SU(N_{f})_{R}$ can exist in QCD only when
@@
When this happens, this colorless fermionic object necessarily carries a baryon number $q_{B}$ satisfying $q_{B}=1\mod N_{f}$. For example, with physical $N_{c}=3$, the above conditions are satisfied by $N_{f}=2$, but are violated by $N_{f}=3$. Therefore, a gapless window at $\theta_{B}=\pi$ described by the free-fermion CFT (S2.1) is at least conceivable for $N_{f}=2$, but is immediately excluded for $N_{f}=3$. Of course, a more complicated free-fermion CFT might simultaneously match the symmetry and the anomaly, and we leave a detailed exploration of such possibilities to the future.
@@
The beta function of the $(2+\epsilon)$D $SU(N_{f})$ PCM is known up to the four-loop order [62, 63, 119] (their coupling $t$ in the Lagrangian is related to our $g^{2}$ through $g^{2}=4t$):
[…]
where $\zeta_{3}\approx 1.202$ is the Riemann zeta function evaluated at $3$. The UV fixed point is thus located at
$g_{*}^{2}\ =\ \frac{4}{N_{f}}\epsilon-\frac{2}{N_{f}}\epsilon^{2}+\frac{1}{2N_{f}}\epsilon^{3}-\left(\frac{1}{3N_{f}}+\frac{3\zeta_{3}}{N_{f}^{3}}\right)\epsilon^{4}+\mathcal{O}(\epsilon^{5})$ (S4.2)
The scaling dimension of the energy operator $\mathcal{O}_{E}$ is related to the critical exponent $\nu$ that governs the divergence of the correlation length as the UV fixed point:
@@
$\begin{split}\Delta_{E}(\epsilon)\ =\ 2+\epsilon-\frac{1}{\nu}\ =\ 2+\epsilon+\beta^{\prime}(g^{2}_{*})\ =\ 2-\frac{1}{2}\epsilon^{2}-\frac{1}{4}\epsilon^{3}-\left(\frac{5}{16}+\frac{9\zeta_{3}}{4N_{f}^{2}}\right)\epsilon^{4}+\mathcal{O}(\epsilon^{5})\,.\end{split}$ (S4.3)
This series appears to be non-alternating. Curiously, $N_{f}$-dependence only shows up from the fourth order.
@@
The anomalous dimension of the vertex operator $\mathcal{O}_{\phi}$ in the $\boldsymbol{N_{f}}\otimes\overline{\boldsymbol{N_{f}}}$ representation of $SU(N_{f})_{L}\times SU(N_{f})_{R}$ is known up to the three-loop order [87] (their coupling $g^{2}$ in the Lagrangian is the same as ours):
$\gamma_{\phi}(g^{2})\ =\ \frac{\left(N_{f}^{2}-1\right)}{4N_{f}}\left(g^{2}+\frac{3N_{f}^{2}}{128}g^{6}\right)+\mathcal{O}\left(g^{8}\right)\,.$ (S4.4)
@@
This $\gamma_{\phi}$ is the anomalous dimension with respect to the engineering dimension 0. Therefore, this $\gamma_{\phi}$ is equivalent to $\frac{1}{2}\zeta$ used in Ref. [61, Section 4.1.3] (see especially their Footnote 35) for the $O(N)$ nonlinear sigma model. Hence the scaling dimension of $\phi$ at the UV fixed point is given by
$\Delta_{\phi}(\epsilon)\ =\ \gamma_{\phi}(g_{*}^{2})\ =\ \left(1-\frac{1}{N_{f}^{2}}\right)\left[\epsilon-\frac{1}{2}\epsilon^{2}+\frac{1}{2}\epsilon^{3}\right]+\mathcal{O}\left(\epsilon^{4}\right)\,.$ (S4.5)
This series appears to be alternating. The factorization of the $N_{f}$-dependence is unlikely to hold from the fourth order.
To extrapolate $\Delta_{E}(\epsilon)$ and $\Delta_{\phi}(\epsilon)$ to $\epsilon=1$, we need to deal with a standard challenge. Since there are no non-trivial fixed points for $\epsilon<0$, Dyson’s argument [37] suggests that the perturbative expansion in $\epsilon$ should be asymptotic. Simply plugging $\epsilon=1$ into a truncated series representation of $\Delta_{E}(\epsilon)$ and $\Delta_{\phi}(\epsilon)$ is thus likely to be a very poor approximation. However, it is natural to expect observables like $\Delta_{E}(\epsilon),\Delta_{\phi}(\epsilon)$ to have resurgent transseries representations around $\epsilon=0$, see e.g. Ref. [4] for a review. The standard way to extrapolate asymptotic series is via Borel summation of an analytic continuation of the Borel transform of the series. Here we will adopt the simplest and most widely used analytic continuation, which is via Padé approximants, which are rational functions that match the series expansion to a given order. We will use this Padé-Borel (PB) resummation approach to estimate the values of $\Delta_{E}(\epsilon)$ and $\Delta_{\phi}(\epsilon)$ at $\epsilon=1$. To estimate the errors on our estimates, we will use combinations of the variation of our results as we vary the choice of Padé approximant together with the non-perturbative ambiguities in the resummations (when such ambiguities appear).
S4.1 Extrapolation of $\Delta_{E}$
Replacing $\epsilon^{n}$ with $\delta^{n}/n!$, we obtain the Borel transform of the truncated $\Delta_{E}(\epsilon)$ series (S4.3):
$\mathcal{B}[\Delta_{E}](\delta)\ =2-\frac{\delta^{2}}{4}-\frac{\delta^{3}}{24}-\left(\frac{5}{16}+\frac{9\zeta_{3}}{4N_{f}^{2}}\right)\frac{\delta^{4}}{24}+\mathcal{O}\left(\delta^{5}\right)\,.$ (S4.6)
If $\Delta_{E}(\epsilon)$ is indeed non-alternating, we would expect $\mathcal{B}[\Delta_{E}](\delta)$ to have a leading singularity on the positive real axis. Such a singularity leads to imaginary ambiguities in the Borel integral that must be canceled by appropriate nonperturbative contributions in a resurgent transseries representation of $\Delta_{E}(\epsilon)$.